OpenStudy (mandre):

I have a question related to Numerical methods which is a Comp Sci module so I figured it belongs here. My question is: whenever a question states that an answer must be rounded/chopped to X digits, is it always expected of me to normalize? I'm in a study group and they all state that it is not necessary unless it's specifically asked for.

OpenStudy (mandre):

I also believe I should normalize after every operation like this: Using 4-digit chopping/truncation: 0.22345 - (23.1864 + 6.02065) / 0.00220239 This becomes: 0.2234 - (23.18 + 6.02) / 0.002202 = 0.2234 - (29.2) / 0.002202 = 0.2234 - 13260.67211625795 = 0.2234 - 13260 = -13259.7766 = -13250 Am I doing this right?